OpenStudy (anonymous):

The Cool Company determines that the supply function for its basic air conditioning unit is S(p) = 40 + 0.008p3 and that its demand function is D(p) = 200 - 0.16p2, where p is the price. Determine the price for which the supply equals the demand

OpenStudy (anonymous):

\[40+0.008p^3=200-0.16p^2\\ 0.008p^3+0.16p^2-160=0\] http://www.wolframalpha.com/input/?i=0.008p%5E3%2B0.16p%5E2-160%3D0

OpenStudy (anonymous):

Ok thanks could you help with this next one ?